# Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles with Reliable Explanations

### v4

@@ C.2 User Study
C.3 Ablation Study
To validate the effect of different components in our methodology, we either vary or remove them and observe their impact on model performance, as measured by F1@Val in Table A1. First, we observe that fine-tuning is the most important component, as its removal causes a dramatic performance drop. This indicates it is necessary to adapt LLMs to this task of fraud detection. External knowledge is very beneficial, and the removal leads to a significant decrease. Meanwhile, other components can provide meaningful benefits, which confirms the effectiveness of our methodology.
@@ C.3 Ablation Study
Prompt of Detecting Fraudulent Articles
@@
Evaluate the relevance of explanation provided for the model’s prediction of the fraudulent articles.
@@
Evaluate the coherence of the explanation provided for the model’s prediction of fraudulent articles. Consider the following criteria:
@@
In the meantime, a set of legitimate papers is provided, each containing highly relevant articles. Use these articles as evidence to assess whether the given paper is potentially fraudulent. Look for indicators such as controversial topics, unsupported claims, or inconsistencies. Below are examples of relevant papers:

### v5

@@ C.3 User Study
C.4 Ablation Study
To validate the effect of different components in our methodology, we either vary or remove them and observe their impact on model performance, as measured by F1@Val in Table A2. First, we observe that fine-tuning is the most important component, as its removal causes a dramatic performance drop. This indicates it is necessary to adapt LLMs to this task of retraction detection. External knowledge is very beneficial, and the removal leads to a significant decrease. Meanwhile, other components can provide meaningful benefits, which confirms the effectiveness of our methodology.
Test Breast Lung Ovarian Colorectal Avg
Vanilla 84.8 88.5 89.2 85.9 69.7 83.6
RAG 66.9 82.7 87.9 85.6 81.8 81.0
Debate 90.7 92.2 94.6 89.9 75.9 88.7
Table A1: Recall (average over 3 seeds) across modes
@@ C.4 Ablation Study
Prompt of Detecting Retracted Articles
@@
Evaluate the relevance of explanation provided for the model’s prediction of the retracted articles.
@@
Evaluate the coherence of the explanation provided for the model’s prediction of retracted articles. Consider the following criteria:
@@
In the meantime, a set of legitimate papers is provided, each containing highly relevant articles. Use these articles as evidence to assess whether the given paper is potentially retracted. Look for indicators such as controversial topics, unsupported claims, or inconsistencies. Below are examples of relevant papers: